# Crystal-field effects in the formation of Wigner-molecule supercrystals in moiré TMD superlattices

### v1

@@ Abstract
For moiré bilayer TMD superlattices, full-configuration-interaction (FCI) calculations are presented that take into account both the intra-moiré-quantum-dot (MQD) charge-carrier Coulombic interactions, as well as the crystal-field effect from the surrounding moiré pockets (inter-moiré-QD interactions). Such FCI calculations enable an effective computational embedding strategy and allow for a complete interpretation of the counterintuitive experimental observations reported recently in the context of moiré TMD superlattices at integer fillings $\nu=2$ and 4. Two novel states of matter are reported: (i) a genuinely quantum-mechanical supercrystal of sliding Wigner molecules (WMs) for unstrained moiré TMD materials (when the crystal field is commensurate with the trilobal symmetry of the confining potential in each embedded MQD) and (ii) a supercrystal of pinned Wigner molecules when strain is involved and the crystal field is incommensurate with the trilobal symmetry of the confining potential in each embedded MQD. The case of $\nu=3$ is an exception, in that both unstrained and strained cases produce a supercrystal of pinned WMs, which is due to the congruence of intrinsic (that of the WM) and external (that of the confining potential of the MQD) $C_{3}$ point-group symmetries. Furthermore, it is shown that the unrestricted Hartree-Fock approach fails to describe the supercrystal of sliding WMs in the unstrained case, providing a qualitative agreement only in the case of a supercrystal of pinned WMs.
I Introduction
Earlier studies have revealed a novel fundamental-physics aspect in artificial twodimensional (2D) nanosystems in the regime of strong interelectronic correlations, namely formation of quantum Wigner molecules (WMs), originally described theoretically [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26] in 2D semiconductor (parabolic, elliptic, and double-well) quantum dors (QDs), as well as in trapped ultracold atoms, and subsequently observed experimentally in GaAs QDs [27, 28, 29, 30], Si/SiGe QDs [31], planar germanium QDs [32] and 1D carbon-nanotubes [33].
@@ I Introduction
Building on the demonstrated formation [34] of WMs in the isolated moiré pockets [most often referred to as moiré quantum dots (MQDs)] at integer fillings, we investigate here the effects arising from embedding such single MQDs in the moiré superlattice structure, which is the actual system that was addressed experimentally [36]. Specifically, this paper adresses the key question whether the experimental observations in the unstrained-lattice case do reflect formation of a collective WM supercrystal, i.e., a novel, more complex insulating phase of matter beyond the well-known generalized Wigner crystal, or the trivial case of a collection of independent MQDs. The present investigations (see below) support formation of a WM supercrystal  [^1] We note that the experimental results for the unstrained case in Ref. [36] were necessarily interpreted with formation of sliding WMs in isolated MQDs (for $N=2$ confined charge carriers, see Refs. [34, 36]; for $N=4$ confined charge carriers, see Ref. [34]), given the fact that the mean-field Hartree-Fock approach (that allows calculations for a large number of charge carriers distributed over several moiré pockets) is unreliable [35, 36] for describing sliding WMs. having sliding WMs as building blocks for $\nu=2$ or $\nu=4$, with the filling $\nu=3$ being an exception having a pinned WM as a building block.
Specifically, this paper focuses on the effects on WM formation in a given MQD resulting from the crystal-field generated by the charge carriers (of same nature) in the surrounding MQDs (moiré pockets) located at the apices of a regular hexagon in the TMD moiré superlattice. This approach is inspired by the well-known crystal-field theory [38, 39] in molecular physics, and is applicable here because the moiré lattice constant $a_{M}$ is much larger than the extent of the Wigner molecule within a single MQD. Two different computational methodologies will be used in this endeavor, namely: (i) the full configuration interaction (FCI) [40, 8, 9, 41, 42, 14, 17, 43, 21, 22, 26, 26, 44, 45] and (ii) the spin-and-space unrestricted Hartree-Fock (sS-UHF) [1, 46, 5, 14, 47].
II Many-body Hamiltonian (including crystal field from surrounding moiré pockets)
Following earlier established literature [48, 49, 50, 34, 36], we approximate the potential of the 2D TMD moiré superlattice that confines the extra charge carriers as
@@ II Many-body Hamiltonian (including crystal field from surrounding moiré pockets)
A brief outline of the FCI and sS-UHF methodologies, used to solve the corresponding many-body Schrödinger equation, is presented in the Appendices. Before proceeding with the computational results, we point out that the arrangement having the equilateral three-point-charge distributions in all the six surrounding moiré pockets [see schematic in Fig. 1(a)] displays an overall $C_{6}$ point-group symmetry, which is commensurate with the trilobal $C_{3}$ symmetry associated with the potential confinement [Eq. (2)] of a single MQD. Below, we will investigate the effect of both such a commensurate crystal field, as well as the effect of an incommensurate crystal field (IcCF) generated by placing the three-point-charge distribution in only one of the six surrounding moiré pockets [see schematic in Fig. 1(c)].
Figure 2: sS-UHF ground-state (with total-spin projection $S_{z}=0$) charge densities for $N=2$ holes in the central MQD (red triangle in the schematics) taking into consideration the crystal-field effect generated by the charge carriers in the surrounding six moiré pockets (see blue hexagon, with radius $a_{M}$, in the schematics). (a) and (c) $\kappa=5$ in agreement with the case of a TMD moiré superlattice in hBN environment. (b) and (d) were calculated, for purpose of comparison, with $\kappa=2$, corresponding to a stronger Coulomb repulsion. In (a) and (b), all six surrounding MQDs (see schematics) were populated with $Q=Ne$ charge carriers. In (c) and (d), only one of the surrounding MQDs (the one to the right, see schematics) has been populated with $Q=Ne$ charge carriers. The charge-carrier distributions in each surrounding MQD have been represented by three point charges $Q/3$ placed at the apices of an equilateral triangle (see schematics), with radius $a_{M}/6$. Remaining parameters: effective mass $m^{*}=0.90m_{e}$, moiré lattice constant $a_{M}=9.8$ nm, depth of a moiré pocket $v_{0}=10.3$ eV, and $\phi=20^{\circ}$. Charge densities are in units of 1/nm2.
III Results for filling factor $\nu=2$
The FCI and sS-UHF charge densities (CDs) for a moiré TMD (e.g., WS2 [36]) superlattice at integer filling factor $\nu=2$ calculated using the many-body Hamiltonian (…) are displayed in Fig. 1 and Fig. 2, respectively. The configuration of the equilateral three-point-charge distributions that mimick the crystal-field effect from the surrounding six moiré pockets is illustrated in the schematics drawn in the upper right corner of each frame; see figure captions for details and parameters. The total charge associated with each equilateral three-point-charge distribution is $Q=2e$. (Note that $Q/e$ can be taken different than $\nu$ in the case of a noninteger filling.)
@@ III Results for filling factor $\nu=2$
Figure 3: Charge densities for $N=4$ holes in the central MQD (red triangle in the schematics) taking into consideration the crystal-field effect generated by the charge carriers in the surrounding six moiré pockets (see blue hexagon, with radius $a_{M}$, in the schematics). (a) and (b) FCI ground-states (with total spin $S=1$ and total-spin projection $S_{z}=0$). (c) and (d) sS-UHF ground-states (with total spin projection $S_{z}=0$). Dielectric constant $\kappa=5$ in agreement with the case of a TMD moiré superlattice in an hBN environment. In (a) and (c), all six surrounding MQDs (see schematics) were populated with $Q=Ne$ charge carriers. In (b) and (d), only one of the surrounding MQDs (the one to the right, see schematics) has been populated with $Q=Ne$ charge carriers. The charge-carrier distributions in each surrounding MQD have been represented by three point charges $Q/3$ placed at the apices of an equilateral triangle (see schematics), with radius $a_{M}/6$. Remaining parameters: effective mass $m^{*}=0.90m_{e}$, moiré lattice constant $a_{M}=9.8$ nm, depth of a moiré pocket $v_{0}=10.3$ eV, and $\phi=20^{\circ}$. Charge densities are in units of 1/nm2.
Figure 4: Charge densities for $N=3$ holes in the central MQD (red triangle in the schematics) taking into consideration the crystal-field effect generated by the charge carriers in the surrounding six moiré pockets (see blue hexagon, with radius $a_{M}$, in the schematics). (a) FCI ground-states (with total spin $S=1/2$ and total-spin projection $S_{z}=1/2$). (b) sS-UHF ground-states (with total spin projection $S_{z}=1/2$). Dielectric constant $\kappa=5$ in agreement with the case of a TMD moiré superlattice in an hBN environment. In (a) and (b), all six surrounding MQDs (see schematics) were populated with $Q=Ne$ charge carriers. The charge-carrier distributions in each surrounding MQD have been represented by three point charges $Q/3$ placed at the apices of an equilateral triangle (see schematics), with radius $a_{M}/6$. Remaining parameters: effective mass $m^{*}=0.90m_{e}$, moiré lattice constant $a_{M}=9.8$ nm, depth of a moiré pocket $v_{0}=10.3$ eV, and $\phi=20^{\circ}$. Charge densities are in units of 1/nm2.
@@ IV Results for filling factor $\nu=4$
The ability of the crystal-field with the commensurate $C_{6}$ symmetry to preserve, and even enhance, the formation of a quantum sliding WM in the central MQD is revealed in an even more spectacular way in the case of the filling factor $\nu=4$, where Ref. [36] has reported within each MQD the observation of CDs with 3 humps ($C_{3}$ symmetry) for the unstrained lattice, but with 4 humps (broken symmetry) for the strained lattice. Our FCI CDs that take into account the crystal-field effect are in agreement with the experiment; see the 3-hump, sliding-WM CD in Fig. 3(a) (commensurate crystal field) and the 4-hump, distorted-pinne-WM CD in Fig. 3(b) (IcCF). We further note that the sS-UHF fails to describe the sliding WM, yielding instead a broken-$C_{3}$-symmetry CD (pinned WM) with 4 well defined humps even for the case of the commensurate crystal field; see Fig. 3(c). In the case of an IcCF, the sS-UHF agrees qualitatively with the FCI CD as it describes a deformed, 4-hump, pinned WM; compare Fig. 3(b) and Fig. 3(d), the charge-carrier localization being more pronounced in the sS-UHF CD.
V Results for filling factor $\nu=3$
@@
Compared to the cases of $\nu=2$ and $\nu=4$, the case of $\nu=3$ is exceptional, a fact that was reported already in Ref. [34] where formation of WMs was investigated for the case of an isolated MQD in the absence of a crystal field from the surrounding moiré pockets. Specifically, the 3-hump pinned WM in the central MQD has a $C_{3}$ symmetry which coincides with that of the confining potential. This congruence of intrinsic and external point-group symmetries prohibits the formation of a sliding WM and yields a strongly pinned WM. The commensurate crystal field does not alter this CD distribution, a fact that is reflected in the FCI calculation [see Fig. 4(a)]. In the $\nu=3$ case, the sS-UHF result is in qualitatively aggreement with the FCI one [see Fig. 4(b)], with the charge-carrier localization, however, being more pronounced in the sS-UHF case. Again, it is remarkable that our FCI results (including the crystal-field effect) for $\nu=3$ are in excellent agreement with the experiment [36].
VI Summary
@@
Taking into consideration (via a computational embedding scheme) both the intra-moiré-QD charge-carrier interactions in the embedded MQD, as well as the crystal-field effect from the surrounding moiré pockets (inter-moiré-QD interactions), and using the full-configuration-interaction methodology, we offered a complete interpretation of the counterintuitive experimental observations reported in Ref. [36] in the context of moiré TMD superlattices at integer fillings $\nu=2$ and 4. In particular, our results demonstrate that these experimental observations reflect the interplay between two novel states of matter; namely, (i) a supercrystal of sliding Wigner molecules  [^2] When the QD confinement possesses circular symmetry, the sliding WMs were earlier referred to as rotating WMs or rotating electron molecules [14, 8]. for unstrained moiré TMD materials (when the crystal field is commensurate with the trilobal symmetry of the confining potential in each MQD) and (ii) a supercrystal of pinned Wigner molecules when strain is involved and the crystal field is non-commensurate with the trilobal symmetry of the confining potential in each MQD. The case of $\nu=3$ is an exception, in that both unstrained and strained cases produce a supercrystal of pinned WMs; this is due to the congruence of intrinsic (that of the WM) and external (that of the confining potential of the MQD) $C_{3}$ point-group symmetries.
Furthermore, we demonstrated that the UHF approach (invoked in Refs. [36, 56]) utterly fails to describe the genuinely quantum-mechanical supercrystal of sliding WMs in the unstrained case, providing a qualitative agreement only in the case of a supercrystal of pinned WMs.
The insights gained in this work through systematic comparisons, employing the UHF method with exact diagonalization (FCI) calculations as benchmarks, provide impetus and pathways for the development of effective strategies for advancing economically feasible and accurate computational methodologies for treating strongly-correlated many-body states in materials, including the electron (hole) Wigner-molecular TMD superlattice systems studied here.
VII ackowledgements
@@
This work has been supported by a grant from the Air Force Office of Scientific Research (AFOSR) under Award No. FA9550-21-1-0198. Calculations were carried out at the GATECH Center for Computational Materials Science.
Appendix A THE CONFIGURATION INTERACTION METHOD

### v2

@@ Abstract
For moiré bilayer TMD superlattices, full-configuration-interaction (FCI) calculations are presented that take into account both the intra-moiré-quantum-dot (MQD) charge-carrier Coulombic interactions, as well as the crystal-field effect from the surrounding moiré pockets (inter-moiré-QD interactions). The effective computational embedding strategy introduced here allows for an FCI methodogy that enables the complete interpretation of the counterintuitive experimental observations reported recently in the context of moiré TMD superlattices at integer fillings $\nu=2$ and 4. Two novel states of matter are reported: (i) a genuinely quantum-mechanical supercrystal of sliding Wigner molecules (WMs) for unstrained moiré TMD materials (when the crystal field is commensurate with the trilobal symmetry of the confining potential in each embedded MQD) and (ii) a supercrystal of pinned Wigner molecules when the crystal field is incommensurate with the trilobal symmetry or straining of the whole material is involved. The case of $\nu=3$ is an exception, in that even the unstrained case is associated with a supercrystal of pinned WMs, which is due to the congruence of intrinsic (that of the WM) and external (that of the confining potential of the MQD) $C_{3}$ point-group symmetries. Furthermore, it is shown that the unrestricted Hartree-Fock approach fails to describe the supercrystal of sliding WMs in the unstrained case, providing a qualitative agreement only in the case of a supercrystal of pinned WMs.
Introduction: Earlier studies have revealed a novel fundamental-physics aspect in artificial twodimensional (2D) nanosystems in the regime of strong interelectronic correlations, namely formation of quantum Wigner molecules (WMs), originally described theoretically [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26] in 2D semiconductor (parabolic, elliptic, and double-well) quantum dors (QDs), as well as in trapped ultracold atoms, and subsequently observed experimentally in GaAs QDs [27, 28, 29, 30], Si/SiGe QDs [31], planar germanium QDs [32] and 1D carbon-nanotubes [33].
@@
Building on the demonstrated formation [34] of WMs in the isolated moiré pockets [most often referred to as moiré quantum dots (MQDs)] at integer fillings, we investigate here the effects arising from embedding such single MQDs in the moiré superlattice structure, which is the actual system that was addressed experimentally [36]. Specifically, this paper adresses the key question whether the experimental observations in the unstrained-lattice case do reflect formation of a collective WM supercrystal, i.e., a novel, more complex insulating phase of matter beyond the well-known generalized Wigner crystal, or the trivial case of a collection of independent MQDs. The present investigations (see below) support formation of a WM supercrystal  [^1] We note that the experimental results for the unstrained case in Ref. [36] were necessarily interpreted with formation of sliding WMs in isolated MQDs (for $N=2$ confined charge carriers, see Refs. [34, 36]; for $N=4$ confined charge carriers, see Ref. [34]), given the fact that the mean-field Hartree-Fock approach (that allows calculations for a large number of charge carriers distributed over several moiré pockets) is unreliable [35, 36, 71, 69, 70] for describing sliding WMs. having sliding WMs as building blocks for $\nu=2$ or $\nu=4$. The filling $\nu=3$ is an exception having a pinned WM as a building block.
(We note that we reserve the term "rotating WM" to describe the azimuthal degree of freedom of WMs forming only in a circularly symmetric confinement [3, 14, 8, 13, 5, 10, 17, 22, 12, 38] (thus exhibiting a good total angular momentum), whereas the TMD moiré pockets are distinguished by an azimuthal $C_{3}$ trilobal symmetry (see discussion of the confining potential below), which does not preserve angular momenta.)
Specifically, this paper focuses  [^2] The main text focuses on charge densities (first-order correlations). Results for conditional probability distributions (second-order correlations) are presented in Appendix A and Appendix G. on the effects on WM formation in a given MQD resulting from the crystal-field generated by the charge carriers (of same nature) in the surrounding MQDs (moiré pockets) located at the apices of a regular hexagon in the TMD moiré superlattice. This approach is inspired by the well-known crystal-field theory [40, 41] in molecular physics, and is applicable here because the moiré lattice constant $a_{M}$ is much larger than the extent of the Wigner molecule within a single MQD. Two different computational methodologies will be used in this endeavor, namely: (i) the full configuration interaction (FCI) [42, 43, 8, 9, 38, 44, 14, 17, 45, 21, 22, 26, 26, 46, 47] and (ii) the spin-and-space unrestricted Hartree-Fock (sS-UHF) [1, 48, 5, 14, 49, 50].
Many-body Hamiltonian (including crystal field from surrounding moiré pockets): Following earlier established literature [51, 52, 53, 34, 36], we approximate the potential of the 2D TMD moiré superlattice that confines the extra charge carriers as
@@
A brief outline of the FCI and sS-UHF methodologies, used to solve the corresponding many-body Schrödinger equation, is presented in Appendices C-E, including Refs. [54, 55].
Before proceeding with the computational results, we point out that the arrangement having the equilateral three-point-charge distributions in all the six surrounding moiré pockets [see schematic in Fig. 1(a)] displays an overall $C_{6}$ point-group symmetry, which is commensurate with the trilobal $C_{3}$ symmetry associated with the potential confinement [Eq. (2)] of a single MQD. Below, we will investigate the effect of both such a commensurate crystal field, as well as the effect of an incommensurate crystal field (IcCF) generated by placing the three-point-charge distribution in only one of the six surrounding moiré pockets [see schematic in Fig. 1(c)].
Figure 2: sS-UHF ground-state (with total-spin projection $S_{z}=0$) charge densities for $N=2$ holes in the central MQD (red triangle in the schematics) taking into consideration the crystal-field effect generated by the charge carriers in the surrounding six moiré pockets (see blue hexagon, with radius $a_{M}$, in the schematics). The information in this figure is organized in the same way as in Fig. 1.
Results for filling factor $\nu=2$: The FCI and sS-UHF charge densities (CDs) for a moiré TMD (e.g., WS2 [36]) superlattice at integer filling factor $\nu=2$ calculated using the many-body Hamiltonian (3) are displayed in Fig. 1 and Fig. 2, respectively. The configuration of the equilateral three-point-charge distributions that mimick the crystal-field effect from the surrounding six moiré pockets is illustrated in the schematics drawn in the upper right corner of each frame; see figure captions for details and parameters. The total charge associated with each equilateral three-point-charge distribution is $Q=2e$. (Note that $Q/e$ can be taken different than $\nu$ in the case of a noninteger filling.)
@@
Figure 3: Charge densities for $N=4$ holes in the central MQD (red triangle in the schematics) taking into consideration the crystal-field effect generated by the charge carriers in the surrounding six moiré pockets (see blue hexagon, with radius $a_{M}$, in the schematics). (a) and (b) FCI ground-states (with total spin $S=1$ and total-spin projection $S_{z}=0$). (c) and (d) sS-UHF ground-states (with total spin projection $S_{z}=0$). $\kappa=5$ in all panels. In (a) and (c), all six surrounding MQDs (see schematics) were populated with $Q=Ne$ charge carriers. In (b) and (d), only one of the surrounding MQDs (the one to the right, see schematics) has been populated with $Q=Ne$ charge carriers. All other parameters and notations are the same as those in Figs. 1 and 2.
@@
Results for filling factor $\nu=4$: The ability of the crystal-field with the commensurate $C_{6}$ symmetry to preserve, and even enhance, the formation of a quantum sliding WM in the central MQD is revealed in an even more spectacular way in the case of the filling factor $\nu=4$, where Ref. [36] has reported within each MQD the observation of CDs with 3 humps ($C_{3}$ symmetry) for the unstrained lattice, but with 4 humps (broken symmetry) for the strained lattice. Our FCI CDs that take into account the crystal-field effect are in agreement with the experiment; see the 3-hump, sliding-WM CD in Fig. 3(a) (commensurate crystal field) and the 4-hump, distorted-pinned-WM CD in Fig. 3(b) (IcCF). We further note that the sS-UHF fails to describe the sliding WM, yielding instead a broken-$C_{3}$-symmetry CD (pinned WM) with 4 well defined humps even for the case of the commensurate crystal field; see Fig. 3(c). In the case of an IcCF, the sS-UHF agrees qualitatively with the FCI CD as it describes a deformed, 4-hump, pinned WM; compare Fig. 3(b) and Fig. 3(d), the charge-carrier localization being more pronounced in the sS-UHF CD.
Figure 4: Charge densities for $N=3$ holes in the central MQD (red triangle in the schematics) taking into consideration the crystal-field effect generated by the charge carriers in the surrounding six moiré pockets (see blue hexagon, with radius $a_{M}$, in the schematics). (a) FCI ground-states (with total spin $S=1/2$ and total-spin projection $S_{z}=1/2$). (b) sS-UHF ground-states (with total spin projection $S_{z}=1/2$). $\kappa=5$ in both panels. All other details are the same as those in Figs. 1-3.
The central role of the overall hexalobular $C_{6}$ symmetry in enhancing the formation of sliding WMs is further demonstrated by our calculations for $N=4$ holes using a point-charge arrangement in the surrounding moiré pockets that is antiparallel to that employed in Fig. 3(a). We find that the effect of such a flip is minimal, with the central charge density retaining the same shape and orientation as in Fig. 3(a); see Fig. 8 in Appendix F.
@@
Results for filling factor $\nu=3$: Compared to the cases of $\nu=2$ and $\nu=4$, the case of $\nu=3$ is exceptional, a fact that was reported already in Ref. [34] where formation of WMs was investigated for the case of an isolated MQD in the absence of a crystal field from the surrounding moiré pockets. Specifically, the 3-hump pinned WM in the central MQD has a $C_{3}$ symmetry which coincides with that of the confining potential. This congruence of intrinsic and external point-group symmetries prohibits the formation of a sliding WM and yields a strongly pinned WM. The commensurate crystal field does not alter this CD distribution, a fact that is reflected in the FCI calculation [see Fig. 4(a)]. In the $\nu=3$ case, the sS-UHF result is in qualitatively agreement with the FCI one [see Fig. 4(b)], with the charge-carrier localization, however, being more pronounced in the sS-UHF case. Again, it is remarkable that our FCI results (including the crystal-field effect) for $\nu=3$ are in excellent agreement with the experiment [36].
Explicit consideration of strain: Heretofore, we focused on the influence of the symmetry of the crystal field on the formation of sliding or pinned WMs. Another factor is the presence of strain in the material. This case is examined in some detail in Appendix B, where Fig. 7 displays the charge densities for $N=2$ holes for values of the strain parameter $\delta=0.01$, 0.03, 0.1, and 0.15; $\delta$ is defined as in Ref. [36]. It is seen that a well defined pinned WM is formed as a function of increasing $\delta$, in close agreement with the experiment. It is remarkable that the effect of the breaking of the $C_{3}$ symmetry in the WM charge density can be resolved for small values of $\delta$, as small as 0.03.
@@
Summary: Taking into consideration (via a computational embedding scheme) both the intra-moiré-QD charge-carrier interactions in the embedded MQD, as well as the crystal-field effect from the surrounding moiré pockets (inter-moiré-QD interactions), and using the full-configuration-interaction methodology, we offered a complete interpretation of the counterintuitive experimental observations reported in Ref. [36] in the context of moiré TMD superlattices at integer fillings $\nu=2$ and 4. In particular, our results demonstrate that these experimental observations reflect the interplay between two novel states of matter; namely, (i) a supercrystal of sliding Wigner molecules  [^3] When the QD confinement possesses circular symmetry, the sliding WMs were earlier referred to as rotating WMs or rotating electron molecules [14, 8]. for unstrained moiré TMD materials (when the crystal field is commensurate with the trilobal symmetry of the confining potential in each MQD) and (ii) a supercrystal of pinned Wigner molecules when strain is involved and the crystal field is non-commensurate with the trilobal symmetry of the confining potential in each MQD. The case of $\nu=3$ is an exception, in that both unstrained and strained cases produce a supercrystal of pinned WMs; this is due to the congruence of intrinsic (that of the WM) and external (that of the confining potential of the MQD) $C_{3}$ point-group symmetries.
[…]
Furthermore, we demonstrated that the UHF approach (invoked in Refs. [36, 62]) fails to describe the genuinely quantum-mechanical supercrystal of sliding WMs in the unstrained case, providing a qualitative agreement only in the case of a supercrystal of pinned WMs  [^4] We remark (following one of the referees’ suggestion) that the comparative methodology that we used here of employing FCI calculations as benchmarks for assessing the adequacy of UHF calculations, could be adopted for assessing advanced multi-configurational (MC) quantum-chemical computational techniques [e.g., MC Self-Consistent Field, CASSCF, or the Complete-Active-Space Self-Consistent Field (CASSCF)], as well as for improving such techniques (e.g., introducing ’dynamic’ correlations via perturbative contributions), thereby extending the range of computationally accessible multi-particle systems.
@@
This work has been supported by a grant from the Air Force Office of Scientific Research (AFOSR) under Contract No. FA9550-21-1-0198. Calculations were carried out at the GATECH Center for Computational Materials Science. We thank the referees of the paper for their constructive comments.
Appendix A CONDITIONAL PROBABILITY DISTRIBUTIONS AND SLIDING WIGNER MOLECULES
The unrestricted Hartree-Fock method employs different space orbitals for the two different spin directions. Thus, the sS-UHF preserves the spin projection, but allows the total-spin and space symmetries (i.e., the $C_{3}$ trilobal symmetry in the potential pockets of a TMD moiré superlattice) to be broken, resulting in the appearance of a pinned WM; see Figs. 2(a), 2(b), and 3(c)  [^5] An analogous discussion in the case of a twodimensional rotational confining potential was presented in Ref. [14] in the context of semiconductor quantum dots. The many-body sS-UHF wave function is a single Slater determinant.
The FCI method reproduces the exact many-body solution by providing a many-body wave function that is a linear superposition of Slater determinants. As shown in this paper, the FCI charge densities preserve the $C_{3}$ trilobal symmetry of the moiré potential pocket, even in the presence of the $C_{6}$ hexalobular TMD crystal field; see Figs. 1(a), 1(b), and 3(a).
The $C_{3}$-symmetry preserving charge densities in Figs. 1(a), 1(b), and 3(a) do not suggest by themselves any relation to WMs. Nevertheless, for these cases, we used the term "sliding WM" because the intrinsic structure of the associated FCI wave functions consists of a superposition of WM configurations that exhibit quantum fluctuations in the azimuthal orientation (thus the term "sliding"). That this is the case is promptly revealed through the inspection of second-order correlations, often referred to as spin-resolved conditional probability distributions (SR-CPDs).
Figure 5: FCI spin-resolved CPDs ($P_{\downarrow\downarrow}$) for the ground state (with total spin $S=1$ and total-spin projection $S_{z}=0$) for $N=4$ holes in the central MQD taking into consideration the crystal-field effect generated by the charge carriers in the surrounding six moiré pockets (see blue hexagon, with radius $a_{M}$, in the schematics). $\kappa=5$ in all four panels. The black arrows indicate the chosen down spin of the fixed hole which is situated at the four different positions ${\bf r}_{0}$ marked by a solid black dot: (a) ${\bf r}_{0}$=(0,-2.09 nm), (b) ${\bf r}_{0}$=(2.4 nm, 0), (c) ${\bf r}_{0}$=(0, 2.61 nm), and ${\bf r}_{0}$=(-2.4 nm, 0). The red arrows indicate that the plotted surfaces reveal the distributions of a second spin-down hole. All other parameters are the same as those in Fig. 3(a), which portrays the corresponding charge density. The CPDs are in units of 1/nm4 and are normalized to 1/nm2.
Figure 6: FCI spin-resolved CPDs ($P_{\uparrow\downarrow}$) for the ground state (with total spin $S=1$ and total-spin projection $S_{z}=0$) for $N=4$ holes in the central MQD taking into consideration the crystal-field effect generated by the charge carriers in the surrounding six moiré pockets (see blue hexagon, with radius $a_{M}$, in the schematics). $\kappa=5$ in all four panels. The black arrows indicate the chosen down spin of the fixed hole which is situated at the same positions ${\bf r}_{0}$ as in Fig. 5. The red arrows indicate that the plotted surfaces reveal the distributions of a second spin-up hole. All other parameters are the same as those in Fig. 3(a), which portrays the corresponding charge density. The CPDs are in units of 1/nm4 and are normalized to 2/nm2.
The spin-resolved CPD gives the spatial probability distribution of finding a second fermion with spin projection $\sigma$ under the condition that a first fermion is located (fixed) at the reference point ${\bf r}_{0}$ with spin projection $\sigma_{0}$; $\sigma$ and $\sigma_{0}$ can be either up (($\uparrow$( ↑) or down ($\downarrow$). The meaning of a space-only CPD is analogous, but without consideration of spin. Mathematically, the spin-resolved CPDs are defined [14, 16] as
$P_{\sigma\sigma_{0}}({\bf r},{\bf r}_{0})=\langle\Phi^{\rm{FCI}}|\sum_{i\neq j}\delta({\bf r}-{\bf r}_{i})\delta({\bf r}_{0}-{\bf r}_{j})\delta_{\sigma\sigma_{i}}\delta_{\sigma_{0}\sigma_{j}}|\Phi^{\rm{FCI}}\rangle.$ (6)
Figure 7: Evolution of the FCI charge densities for $N=2$ holes as a function of the strain parameter $\delta$, whose value is displayed in each frame. The evolution from three to two peaks can be seen in the contour plots in the $x-y$ plane, as well as in the 3D surfaces. Remaining parameters: effective mass $m^{*}=0.90m_{e}$, moiré lattice constant $a_{M}=9.8$ nm, depth of a moiré potential pocket $v_{0}=10.3$ eV, $\phi=20^{\circ}$, and $\kappa=5.0$. Charge densities are in units of 1/nm2.
In Figs. 5 and 6, we display SR-CPDs for the ground state of the 4-hole sliding WM whose charge density was displayed in Fig. 3(a). Two different families of SR-CPDs are considered, namely: (1) the $P_{\downarrow\downarrow}$ family, which can be described as "fix a hole with spin down and look for another hole with spin down, as well" (see Fig. 5), and (2) the $P_{\uparrow\downarrow}$ family, which can be described as "fix a hole with spin down and look for another hole with spin up" (see Fig. 6). The remaining two SR-CPD families, $P_{\uparrow\uparrow}$ and $P_{\downarrow\uparrow}$, can be simply obtained from those plotted in Figs. 5 and 6 by flipping the spins.
These four families of SR-CPDs reveal an intrinsic structure for the sliding WM which consists of four localized holes, two with up spins and the other two with downs spins in an alternating configuration. Moreover, following similar reasoning as in Ref. [45], one can deduce that the intrinsic spin eigenfunction associated with the ground-state sliding WM of Fig. 3(a) is given by the expression
$\chi^{S=1}_{S_{z}=0}=(|\uparrow\downarrow\uparrow\downarrow\rangle-|\downarrow\uparrow\downarrow\uparrow\rangle)/\sqrt{2}.$ (7)
Unlike the case of a circular confinement [3, 10, 14], the CPD surfaces in Figs. 5 and 6 vary with the azimuthal angle. As mentioned above, this is associated with the fact that the corresponding charge density in Fig. 3(a) has a trilobally deformed doughnut shape.
For completeness, in Fig. 9 (see Appendix G), we display SR-CPDs for the singlet ground state of the 2-hole sliding WM whose charge density was displayed in Fig. 1(a).
Finally, we comment briefly on the approaches that bridge the gap between the sS-UHF and FCI solutions. In the case of a circular confinement, we showed earlier [5, 14] that this can be achieved by restoring the broken symmetry of the sS-UHF solutions via projection techniques. This is equivalent to carrying out a mixing (superposition) of the equivalent (merely rotated) sS-UHF solutions over all the azimuthal angles using coefficients supplied by the projection methodology. In the case of a moiré quantum dot considered here, the mean-field components entering in the mixing are not equivalent in all the azimuthal directions, and thus the Griffin-Hill-Wheeler generator coordinate method [14, 65, 66, 67], which is broader in scope than the symmetry-restoration approach, must be employed [68].
Appendix B EXPLICIT CONSIDERATION OF STRAIN
In Fig. 7, we display FCI CDs for $N=2$ holes as a function of the strain parameter $\delta$. The remaining parameters are the same as for the unstrained ($\delta=0$) CD displayed in Fig. 1(a). The parameter $\delta$ is defined as in Ref. [36], namely, lengths along the $x$-axis are extended by a factor $(1+\delta/2)$, whilst lengths along the $y$-axis are contracted by a factor $(1-\delta/2)$.
Appendix C THE CONFIGURATION INTERACTION METHOD